Find the distance of the point of intersection of the line
x-2/1=y+3/-1=z/3
and the plane 2x-3y+4z+4=0 from the origin.
1
Expert's answer
2019-07-15T09:35:29-0400
"\frac{x-2}{1}=\frac{y+3}{-1}=\frac{z}{3}=t\\
\begin{cases}
x=t+2\\
y=-t-3\\
z=3t
\end{cases}"

then


"2x-3y+4z+4=\\
=2t+4+3t+9+12t+4=17t+17=0\\
\implies t=-1\\
\begin{cases}
x=1\\
y=-2\\
z=-3
\end{cases}"

and the distance from the origin:

"r=\sqrt{(1)^2+(-2)^2+(-3)^2} = \sqrt{14}"